# The 4DLO and other tubing models of $S^{3}$ symmetry

### v1

@@ Abstract
By multiplying on both left and right, we can fix points, and therefore have symmetry about them. By writing $\pm[A\times B]$ Conway and Smith mean the group of actions of the ${\bf{x}}\to{\bf{a}}{\bf{x}}{\bf{b}}$ where $\bf{a}$} and $\bf{b}$} map into 3d rotation groups $A$ and $B$. Figure 7 shows models of symmetry $\pm[C_{2}\times C_{11}]$ and $\pm[T\times C_{6}]$, where $C_{n}$ is the cyclic group of order $n$. That notation soon has many amendments — we rely on the detailed geometric discussion in  [11], especially their Sections 8.4 and 8.6 on the tesseract and 24-cell. We’ll give Conway and Smith’s notation, referring the reader to the concordances and extensions in [3, 4, 11].
As there is cubic symmetry around each point in $V_{24}\cup V_{24}^{\prime}$, this compound of two 24-cells has rotational symmetry $\pm[O\times O]$. Applying any ${\bf{x}}\to{\bf{r}}{\bf{x}}{\bf{p}}=({\bf{r}}{\bf{p}})({\bf{p}}^{-1}{\bf{x}}{\bf{p}})$ rotates $\bf{1}$} about the $\bf{p}$} axis (a rotation fixing $\bf{1}$} that is a symmetry in $O$) and then by left multiplication moves everything taking $\bf{1}$} to ${\bf{r}}{\bf{p}}^{-1}$, another point of $V_{24}\cup V_{24}^{\prime}$, also in $O^{*}$. The model at bottom left of Figure 5 also has vertices $V_{24}\cup V_{24}^{\prime}$, and has this same symmetry, but it is not two dual 24-cells.
Allowing mirror reflections by taking symmetries of the form ${\bf{x}}\to{\bf{p}}\overline{{\bf{x}}}{\bf{r}}$ where $\overline{{\bf{x}}}=2{\text{re}}({\bf{x}})-{\bf{x}}$, we extend these to the full symmetry of this dual pair, named $\pm[O\times O]\cdot 2$. Henceforth we’ll name only the rotational symmetries.
@@
With half of the symmetry of a dual pair, 4DLO shows the full symmetry of a single 24-cell in Figure 2(a), denoted $\pm\frac{1}{2}[O\times O]$. We’ve seen that the vertices $V_{24}$ of a vertex-down 24-cell sort into $V_{8}$, $V_{16}^{+}$ and $V_{16}^{-}$. Any point corresponding to 3-fold rotation, so it happens, is a symmetry that permutes these sets. For example, multiplying by $\omega:=\frac{1}{2}({\bf{i}}+{\bf{j}}+{\bf{k}}+{\bf{1}})$ moves ${\bf{1}}\in V_{8}$ to $\omega\in V_{16}^{+}$ to $\omega^{2}\in V_{16}^{-}$ to ${\bf{1}}\in V_{8}$, etc. We can put a direction on the edges of a 24-cell in this way, shown in Figures 9 (left) and 2 (b) with type (we think) $\pm[T\times T]$.
@@
Because $V_{8}$ is the vertices of a 16-cell, its images $V_{16}^{+}$ and $V_{16}^{-}$ are too, explaining the symmetry in the compound of three 16-cells, shown in Figures 5 and 6, with vertices $V_{24}$. Equivalently, the 16-cell, as well as its dual, the tesseract, has a third of the symmetry of a 24-cell, which Conway and Smith list as $\pm\frac{1}{6}[O\times O]$. Thus too do we find the compound of three tesseracts sharing vertices with a (dual) 24-cell, shown cell-down in Figure 4, or vertex down in 4DLO at left in Figure 1. There is a thus a compound of two compounds of three tesseracts, with the full symmetry $\pm[O\times O]$.
[…]
Figure 9: (left) From the interactive javascript model at chaimgoodmanstrauss.com/4DLO
(right) A prototype lighted 16-cell (Fayetteville, Ark., 2019)
@@
Inspired by Leo Villareal’s Buckyball and Hive: Bleeker Street, we turned to moving lights to illustrate these symmetries, with a first project shown at right in Figure 9. The 4DLO was much more ambitious, consisting of about 300 feet of 1 inch outer diameter polyethylene tubing (with its beautiful capacity for radiating soft light), and 14,000 WS2811b LED lights diffused by strips of packing foam. These are controlled very simply – each light on each strand receives a string of bytes, taking three for its R, G, and B and passing the rest along. The 96 tubes of the model were organized into four quadrants, each with seven separate strands, powered at each node they passed through. A Teensy 4.1 has more than enough processing power and pins for this task, especially when decked out with a full suite of products from its producer PJRC Electronic Products: Four Octos, one for each quarter for the data lines, their Audio Shield for sound, and extra on-board memory. Each individual LED was associated with a particular location in 4D space, measured along the arc of an edge of the 24-cell; at each moment of the program, an edge would be associated with a coloring function, and these would be assigned by using whatever symmetry we wished to highlight. The javascript program at https://chaimgoodmanstrauss.com/4DLO/ allowed us to prototype the geometry and symmetry — you can also rotate the 24-cell into different positions. We used a bespoke scripting language for the sequencing, which we developed making extensive use of Claude.ai (along the way, learning much about the potential and limitations of today’s LLMs). All of our code is at the Github repository github.com/chaimgoodmanstrauss/4DLO.git. Needless to say, we made many mistakes, and learned a lot, much too much to fit here. (We printed four sets of nodes before getting it right.) Give us a call if you would like to know more.
@@ Acknowledgements
This work has been supported by the Gathering for Gardner and the National Museum of Mathematics; and by the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université, Labex Carmin ANR-10-LABX-59-01’); we are especially grateful to unnamed donors, to Cindy Lawrence, and to the MoMath staff for helping bringing 4DLO to life. The first author expresses deep appreciation for the second author’s ability to make these things real.

### v2

@@ 3  Quaternionic groups as the symmetries of $S^{3}$
By writing $\pm[A\times B]$ Conway and Smith [3] mean the group of actions of the form ${\bf{x}}\mapsto{\bf{a}}{\bf{x}}{\bf{b}}$ where the quaternions $\bf{a}$} and $\bf{b}$} correspond to rotations that lie in symmetry groups $A$ and $B$. Figure 9 shows models of symmetry $\pm[C_{2}\times C_{11}]$ and $\pm[T\times C_{6}]$, where $C_{n}$ is the cyclic group of order $n$. That notation soon has many amendments — we rely on the detailed geometric discussion in  [11], especially their Sections 8.4 and 8.6 on the tesseract and 24-cell. We’ll use Conway and Smith’s notation, referring the reader to the concordances and extensions in [3, 4, 11].
@@
As there is cubic symmetry around each point in $V_{24}\cup V_{24}^{\prime}$, this compound of two 24-cells has rotational symmetry $\pm[O\times O]$. Applying any ${\bf{x}}\mapsto{\bf{r}}{\bf{x}}{\bf{p}}=({\bf{r}}{\bf{p}})({\bf{p}}^{-1}{\bf{x}}{\bf{p}})$ rotates $\bf{1}$} about the $\bf{p}$} axis (a rotation fixing $\bf{1}$} that is a symmetry in $O$) and then by left multiplication moves everything taking $\bf{1}$} to ${\bf{r}}{\bf{p}}$, another point of $V_{24}\cup V_{24}^{\prime}$, also in $O^{*}$. The model at left in Figure 6 also has vertices $V_{24}\cup V_{24}^{\prime}$, and has this same symmetry, but it is not two dual 24-cells. Allowing mirror reflections by taking symmetries of the form ${\bf{x}}\mapsto{\bf{p}}\overline{{\bf{x}}}{\bf{r}}$ where $\overline{{\bf{x}}}=2{\text{Re}}({\bf{x}})-{\bf{x}}$, we extend these to the full symmetry of this dual pair, named $\pm[O\times O]\cdot 2$. Henceforth we’ll name only the rotational symmetries. In Figure 2(a) 4DLO shows the full symmetry of a single 24-cell denoted $\pm\frac{1}{2}[O\times O]$, with half the symmetry of a dual pair. We’ve seen that the vertices $V_{24}$ of a vertex-down 24-cell sort into the sets $V_{8}$, $V_{16}^{+}$ and $V_{16}^{-}$. Any point corresponding to 3-fold rotation, so it happens, is a symmetry that permutes these sets. For example, multiplying by ${\bf{w}}:=\frac{1}{2}({\bf{i}}+{\bf{j}}+{\bf{k}}+{\bf{1}})$ moves ${\bf{1}}\in V_{8}$ to ${\bf{w}}\in V_{16}^{+}$, to ${\bf{w}}^{2}\in V_{16}^{-}$, to ${\bf{1}}\in V_{8}$, to ${\bf{w}}\in V_{16}^{+}$, etc. We can put a direction on the edges of a 24-cell in this way, with type (we think) $\pm[T\times T]$, shown at right in Figure 4 and in Figure 2(b).
@@
Because $V_{8}$ is the vertex set of a 16-cell, its images $V_{16}^{+}$ and $V_{16}^{-}$ are too, explaining the symmetry in the compound of three 16-cells with vertices $V_{24}$, and in the self-dual compound of two of those compounds, shown in Figures 6 and 7. Equivalently, the 16-cell, as well as its dual, the tesseract, has a third of the symmetry of a 24-cell, which Conway and Smith list as $\pm\frac{1}{6}[O\times O]$. Dually there is compound of three tesseracts sharing vertices with a (dual) 24-cell, shown cell-down in Figure 5, or vertex down in 4DLO at left in Figure 1, and a compound of two compounds of three tesseracts, with the full symmetry $\pm[O\times O]$.
4  The Four-Dimensional Light Orchestra, a.k.a. 4DLO
@@
Inspired by Leo Villareal’s Buckyball and Hive: Bleeker Street, we turned to moving lights to illustrate these symmetries. A first project is shown in the center in Figure 3. The 4DLO was much more ambitious, consisting of about 300 feet of 1 inch outer diameter polyethylene tubing (with its beautiful capacity for radiating soft light), and 14,000 WS2811b LED lights diffused by strips of packing foam. These are controlled very simply – each light on each strand receives a string of bytes, taking three for its R, G, and B values and passing along the rest. The 96 tubes of the model were organized into four quadrants, each with seven separate strands, powered at each node they passed through. A Teensy 4.1 microcontroller has more than enough processing power and pins for this task, especially when decked out with a full suite of accessories from its producer PJRC Electronic Products: four Octos, one for each quarter for the data lines, an Audio Shield for sound, and extra on-board memory. Each individual LED was associated with a particular location in 4D space, measured along the arc of an edge of the 24-cell; at each moment of the program, an edge would be associated with a coloring function, and these would be assigned by using whatever symmetry we wished to highlight. The javascript program at https://chaimgoodmanstrauss.com/4DLO/ allowed us to prototype the geometry and symmetry — there you can also rotate the 24-cell into different positions. We used a bespoke scripting language for the sequencing, which we developed making extensive use of Claude.ai (along the way, learning much about the potential and limitations of today’s LLMs). All of our code is at the Github repository github.com/chaimgoodmanstrauss/4DLO.git. Needless to say, we made many mistakes, and learned a lot, much too much to fit here. (We printed four complete sets of nodes before getting it right.) Give us a call if you would like to know more.
@@ Acknowledgements
We thank the referees for improving this paper with many helpful comments. Figure 5 courtesy (left) National Museum of Mathematics (MoMath), (right) Gathering For Gardner Foundation (G4G), Stephen Inglima. This work has been supported by G4G; MoMath; Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université, Labex Carmin ANR-10-LABX-59-01’); and Universidad Nacional Autónoma de México. We are especially grateful to unnamed donors, to Cindy Lawrence, and to the MoMath staff for helping bringing 4DLO to life. The first author expresses deep appreciation for the second author’s ability to make these things real.